OpenStudy (anonymous):

To test the effectiveness of a gunpowder mixture, 1 gram was exploded under controlled STP conditions, and the reaction chamber was found to expand to 310 cm3. What volume would these gases occupy at the temperature produced by an uncontrolled explosion (2,200°C) and 2.1 atmospheres of pressure?

OpenStudy (abmon98):

Under Standard Conditions we have a temperature of 298K and a pressure of 1*10^5Pa Volume is measured in m^3 so if we convert 310cm^3 into m^3 we get 0.00031m^3 Pressure changed as stated in your question in 2.1*10^5 and a temperature of (2200+273)Kelvin P1/P2=V2/V1=T1/T2 Use either of these equation